# CaT-Bench: Benchmarking Language Model Understanding of Causal and Temporal Dependencies in Plans

### v1

@@ Abstract
Understanding the abilities of LLMs to reason about natural language plans, such as instructional text and recipes, is critical to reliably using them in decision-making systems. A fundamental aspect of plans is the temporal order in which their steps needs to be executed, which reflects the underlying causal dependencies between them. We introduce CaT-Bench, a benchmark of Step Order Prediction questions, which test whether a step must necessarily occur before or after another in cooking recipe plans. We use this to evaluate how well frontier LLMs understand causal and temporal dependencies. We find that SOTA LLMs are underwhelming (best zero-shot is only $0.59$ in F1), and are biased towards predicting dependence more often, perhaps relying on temporal order of steps as a heuristic. While prompting for explanations and using few-shot examples improve performance, the best F1 result is only $0.73$. Further, human evaluation of explanations along with answer correctness show that, on average, humans do not agree with model reasoning. Surprisingly, we also find that explaining after answering leads to better performance than normal chain-of-thought prompting, and LLM answers are not consistent across questions about the same step pairs. Overall, results show that LLMs’ ability to detect dependence between steps has significant room for improvement.
@@ 1 Introduction
Using notions of consistency to evaluate their robustness, we also show that almost all out-of-the-box LLMs are largely unreliable. Few-shot prompting with retrieved exemplars improves performance and consistency ($0.49\rightarrow 0.68$ F1 for gpt-4o). Explanation-based generation offers another route to improve model performance and reliability on reasoning tasks Camburu et al. (2018); Rajani et al. (2019); Kumar and Talukdar (2020). Prompting LLMs to explain their decisions also improves performance on CaT-Bench ($0.49\rightarrow 0.7$ F1). Despite these gains, there is still a large room for improvement in identifying step dependencies. When also considering the quality of the explanations, the average human ratings for satisfactory answers from SOTA LLM’s is only $\sim$3 (out of 5). Further, contrary to prior findings, using chain-of-thought prompting (CoT), i.e., reasoning before answering Wei et al. (2022b), performs worse than answering first and then explaining it, indicating inconsistencies in model reasoning.
@@ 2 Related Work
Plan understanding tasks involve multiple aspects such as entity states Bosselut et al. (2018); Henaff et al. (2017), linking actions Pareti et al. (2014); Lin et al. (2020); Donatelli et al. (2021), next event prediction Nguyen et al. (2017); Zellers et al. (2019); Zhang et al. (2020a) and more. OpenPI Tandon et al. (2020); Zhang et al. (2024a) enables entity tracking in how-to procedures. Zhang et al. (2020b) formalize several multiple-choice tasks related to step- and goal- relations in procedures. Kiddon et al. (2015) explore predicting dependencies in cooking recipes and related tasks. Similar work has been done on identifying dependencies in multimodal instructions with images and text Pan et al. (2020); Wu et al. (2024). PizzaCommonsense Diallo et al. (2024) is a dataset for learning commonsense about intermediate and implicit steps for cooking recipes, and contains explicit input/output pairs of each action with fine-grained annotations. Choice-75 Hou et al. (2023) aims to study decision branching in plans by generating user scenarios and choices for various steps within the plan. CREPE Zhang et al. (2023) measures how well LLMs understand the comparative likelihood of two events occurring in a procedure. There are a variety of datasets evaluating different aspects of plans, but there is a lack of one that clearly studies the prediction and explanation of temporal ordering constraints on the steps of an instructional plan.
@@ 3 CaT-Bench
We build on this idea to create CaT-Bench, a new dataset of causal dependency questions defined over cooking recipes. Specifically, we make use of the Recipe Flow Graph Corpus Yamakata et al. (2020) containing 300 English language cooking recipes annotated with substep procedure dependencies. For each recipe, this dataset provides a directed acyclic graph (DAG), in which the nodes are steps and directed edges indicate the temporal edge between those steps. If the nodes corresponding to two steps are not connected by a directed path, then they can be performed in any order (with respect to themselves) without changing the recipe result. In other words, two steps are temporally dependent if and only if there is a directed edge from one to the other, and independent otherwise.
For all ordered pairs of steps $(i,j)$ in a plan, we create two binary (yes/no) questions: (i) Must $step_{i}$ happen before $step_{j}$? (ii) Must $step_{j}$ happen after $step_{i}$?. These questions primarily test for for precondition relations (e.g. first question in Figure 2), and the ability to understand effects of steps and how they relate to sub-goals or overall goals of the plan (e.g second question in Figure 2). We pool all such questions from dependent pairs of steps (i.e, the steps where there is a directed path from one step’s node to the other in the recipe DAG) into Dep, and the rest into NonDep  [^1] Note that the answers to all the questions in the Dep set are ‘yes’, and the answers to NonDep questions are ‘no’..
@@
CaT-Bench enables two tasks. Step Order Prediction  elicits binary judgments about dependencies between pairs of steps in a plan, and performance on this task can be evaluated automatically. Step Order Explanation  requires models to provide explanations for its judgment about step dependencies. This involves understanding causal relationships and expressing relevant knowledge about actions in the steps being asked about. Since this is a free-form generation task, these explanations require human evaluation. Note that CaT-Bench does not contain gold, human-written explanations, and we advocate for reference-free human evaluation since there can be multiple valid explanations.
@@ 4.1 Models
We evaluate gpt-4-turbo, gpt-3.5-turbo, gpt-4o, gemini-1.0-pro, gemini-1.5-pro, gemini-1.5-flash and Llama3-8B. These represent a diverse set from different model families and sizes. We evaluate them primarily in zero-shot prompting modes. We consider two settings: (i) generating only an answer (A), and (ii) generating an explanation along with the answer (A + E). The latter represents answering the question and then generating an explanation for it. We also analyze few-shot results for the answer-only (A) setting, and evaluate generic CoT (E + A) prompting. More details about each model can be found in Appendix A and the prompts used in Appendix D.
@@ 4.2 How Good Are Model Predictions?
Table 1 presents the performance of all the models in different settings on Step Order Prediction. We present per-class (Dep and NonDep) precision, recall and F1 score as well as macro average metrics on the class balanced CaT-Bench. We make three main observations:
@@
In the zero-shot answer-only setting (A), gpt-3.5-turbo records the highest F1 score overall of 0.59. gpt-4-turbo is close behind with 0.58, but surprisingly gpt-4o, the most recent frontier model, fares significantly worse at 0.49 F1. All three Gemini models (gpt-3.5-turbo, gemini-1.0-pro, and gemini-1.5-flash) also only manage an F1 of around 0.55. Llama3-8B also fares poorly with an F1 of 0.49. Most models are comparable or barely better than a random baseline F1 of 0.5 on this balanced dataset showing that they are not able to directly answer the dependence question.
@@
Results for adding explanations to answers is shown in the (A + E) rows in Table 1. Five of the seven models, gpt-3.5-turbo and gemini-1.5-flash being the exceptions, have higher performance when also generating explanations. The biggest improvement in F1 is seen in gpt-4o (+0.21). With explanations, the best result is the 0.73 F1 when using gemini-1.5-pro. While this is substantially better than a random baseline, there is still significant room for improvement.
@@
As noted above, using explanations improves the overall performance, translating to more balanced precision/recall values on Dep than when predicting answers alone. Since the bias towards Dep necessarily means bias against NonDep, reduction in bias towards Dep also translates to a more balanced performance on both Dep and NonDep sets. However, even with explanations, the bias towards predicting dependence still remains to some extent for all models. Explanations improve gpt-4o performance the most (+0.36) on NonDep questions. They do not help smaller models (Llama3-8B) identify dependencies better.
@@ 4.3 How Good Are Model Explanations?
Table 2 presents the quality of model generated explanations as judged by human annotators. As expected, larger models are clearly better than the much smaller Llama3-8B on all metrics. There is very little difference between the frontier models, gpt-4o, gpt-4-turbo and gemini-1.5-pro. Avg performance indicates that there is significant room for improvement in the quality of model explanations. On ModAvg, we see that even the best model performance is below 3 (‘neither agree nor disagree’ with model explanation). By this metric, gemini-1.5-pro explanations are worse than GPT-4 even though it generates more correct answers. The difference between Avg and ModAvg indicates models are capable of generating convincing explanations for their wrong answers. They produce explanations which justify the opposite of their answer a significant number of times. In fact, Llama3-8B does so almost half the time. These results show that models have a lot of room for improvement in their ability and reliability to reason about step dependencies in plans.
@@ 5.1 Robustness of Models
We use two measure of consistency to quantify the robustness of the models.
@@
For a pair of steps $(step_{i},step_{j})$, the answer to must $step_{i}$ happen before $step_{j}$ should be the same as the answer to must $step_{j}$ happen after $step_{i}$ regardless of question type. As described in subsection 3.1, we measure this notion of consistency through TC. We make two main observations: (i) Even the most consistent models gpt-4o and gemini-1.5-pro change their answers to the before and after versions of questions in 20+of the cases. The rest are far more inconsistent with gemini-1.5-flash changing its answers for more than 55% of the questions; (ii) Surprisingly, adding explanations reduces answer consistency for most models, with gemini-1.5-pro (+24%) and gpt-4-turbo (+14%) being the only exceptions showing improved consistency upon generating explanations.
@@ 5.2 Chain-of-Thought Struggles
In the experiments thus far, we have asked models to generate explanations for their answers. This can be seen as a answer-then-explain (A + E) approach. In contrast, the standard Chain-of-Thought (CoT) prompting strategy Wei et al. (2022b) can be seen as an explain-then-answer approach (E + A), where we ask the model to generate reasoning or explanation that leads to its answer. In practice, this step-by-step reasoning can be seen as allowing the use of intermediate decoding tokens (like a scratchpad) in service to coming up with a possibly more accurate final answer for many tasks. Table 4 compares performance of CoT prompting [^3] We tried multiple CoT prompts but it had little effect on performance. (E + A) to first predicting the answer and then explaining it (A + E) and simply providing an answer (A), all in the zero-shot setting when using gpt-4o. [^4] We use gpt-4o instead of gemini-1.5-pro due to rate limits on the latter.
@@
Further manual analysis of the explanations shows that (E + A) generates more vacuous explanations (e.g. just stating that a pair of steps are independent without digging deeper into what makes them independent). Additionally, (E + A) makes errors by misunderstanding entity states of a step. Unsurprisingly, both settings suffer from some hallucinations, sometimes missing the content of the steps being asked about. These results are further indicators of brittleness and inconsistencies in models’ reasoning about step dependencies.
@@ 5.3 Effect of Improved Prompting Techniques
We use in-context learning Wei et al. (2022a) with examples selected from the balanced training set using the BM25 Robertson and Zaragoza (2009) algorithm. We use k=5 exemplars and dynamically retrieve exemplars from a held-out set that are closest to the test instance. As expected, few-shot prompting improves binary prediction performance a lot (+0.19). In fact, few-shot performance is almost as good as predicting then explaining with gpt-4o.
@@ 6 Conclusion
Understanding plans requires reasoning about its different aspects such as preconditions and effects. This paper introduces CaT-Bench, a new benchmark to evaluate the causal and temporal reasoning abilities about plans. Despite the remarkable strength of current SOTA LLMs, we find that none of them are very good at understanding whether one step in a plan must precede (or succeed) another. Particularly, they are much worse at knowing when there is not a dependency between steps. We also find that LLM predictions are not robust as measured by two metrics of consistency. Prompting LLMs to provide an answer and then to explain it improves performance significantly, and is even better than chain-of-thought (reasoning followed by answering). Human evaluation of these explanations shows that models have a long way to go at understanding dependencies. Our results highlight the current deficiencies of LLMs in plan-based reasoning, and there is need for improvement before they can be reliably used in the many applications requiring such reasoning.
@@ Acknowledgements
This material is based on research that is supported in part by the Air Force Research Laboratory (AFRL), DARPA, for the KAIROS program under agreement number FA8750-19-2-1003 and in part by the National Science Foundation under the award IIS #2007290. This material is also based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0007.

### v3

@@ Abstract
Understanding the abilities of LLMs to reason about natural language plans, such as instructional text and recipes, is critical to reliably using them in decision-making systems. A fundamental aspect of plans is the temporal order in which their steps need to be executed, which reflects the underlying causal dependencies between them. We introduce CaT-Bench, a benchmark of Step Order Prediction questions, which test whether a step must necessarily occur before or after another in cooking recipe plans. We use this to evaluate how well frontier LLMs understand causal and temporal dependencies. We find that SOTA LLMs are underwhelming (best zero-shot is only $0.59$ in F1), and are biased towards predicting dependence more often, perhaps relying on temporal order of steps as a heuristic. While prompting for explanations and using few-shot examples improve performance, the best F1 result is only $0.73$. Further, human evaluation of explanations along with answer correctness show that, on average, humans do not agree with model reasoning. Surprisingly, we also find that explaining after answering leads to better performance than normal chain-of-thought prompting, and LLM answers are not consistent across questions about the same step pairs. Overall, results show that LLMs’ ability to detect dependence between steps has significant room for improvement.
@@ 1 Introduction
Using notions of consistency to evaluate their robustness, we also show that almost all out-of-the-box LLMs are largely unreliable. Few-shot prompting with retrieved exemplars improves performance and consistency ($0.49\rightarrow 0.68$ F1 for gpt-4o). Explanation-based generation offers another route to improve model performance and reliability on reasoning tasks Camburu et al. (2018); Rajani et al. (2019); Kumar and Talukdar (2020). Prompting LLMs to explain their decisions also improves performance on CaT-Bench ($0.49\rightarrow 0.7$ F1). Despite these gains, there is still a large room for improvement in identifying step dependencies. When also considering the quality of the explanations, the average human ratings for satisfactory answers from SOTA LLM’s is only $\sim$3 (out of 5). Further, contrary to prior findings, using chain-of-thought prompting (CoT), i.e., reasoning before answering Wei et al. (2022b), performs worse than answering first and then explaining it, indicating inconsistencies in model reasoning. [^1] CaT-Bench is available at https://huggingface.co/datasets/vanyacohen/CaT-Bench and the code is at https://github.com/StonyBrookNLP/CaT-Bench.
@@ 2 Related Work
Plan understanding tasks involve multiple aspects such as tracking entity states Bosselut et al. (2018); Henaff et al. (2017), linking actions Pareti et al. (2014); Lin et al. (2020); Donatelli et al. (2021), next event prediction Nguyen et al. (2017); Zellers et al. (2019); Zhang et al. (2020a) and more. OpenPI Tandon et al. (2020); Zhang et al. (2024a) enables entity tracking in how-to procedures. ProPara Dalvi et al. (2018) focuses on describing and understanding scientific processes. XPAD Dalvi et al. (2019) extend ProPara by adding the new task of explaining actions by predicting their dependencies. Zhang et al. (2020b) formalize several multiple-choice tasks related to step- and goal- relations in procedures. Kiddon et al. (2015) explore predicting dependencies in cooking recipes and related tasks. Similar work has been done on identifying dependencies in multimodal instructions with images and text Pan et al. (2020); Wu et al. (2024). PizzaCommonsense Diallo et al. (2024) is a dataset for learning commonsense about intermediate and implicit steps for cooking recipes, and contains explicit input/output pairs of each action with fine-grained annotations. Choice-75 Hou et al. (2023) aims to study decision branching in plans by generating user scenarios and choices for various steps within the plan. CREPE Zhang et al. (2023) measures how well LLMs understand the comparative likelihood of two events occurring in a procedure. There are a variety of datasets evaluating different aspects of plans, but there is a lack of one that clearly studies the prediction and explanation of temporal ordering constraints on the steps of an instructional plan.
@@ 3 CaT-Bench
We build on this idea to create CaT-Bench, a new dataset of causal dependency questions defined over cooking recipes. Specifically, we make use of the Recipe Flow Graph Corpus Yamakata et al. (2020) containing 300 English language cooking recipes annotated with substep procedure dependencies. For each recipe, this dataset provides a directed acyclic graph (DAG), in which the nodes are steps and directed edges indicate the temporal edge between those steps. If the nodes corresponding to two steps are not connected by a directed path, then they can be performed in any order (with respect to themselves) without changing the recipe result. In other words, two steps are temporally dependent if and only if there is a directed path from one to the other, and independent otherwise.
For all ordered pairs of steps $(i,j)$ in a plan, we create two binary (yes/no) questions: (i) Must $step_{i}$ happen before $step_{j}$? (ii) Must $step_{j}$ happen after $step_{i}$?. These questions primarily test for for precondition relations (e.g. first question in Figure 2), and the ability to understand effects of steps and how they relate to sub-goals or overall goals of the plan (e.g second question in Figure 2). We pool all such questions from dependent pairs of steps (i.e, the steps where there is a directed path from one step’s node to the other in the recipe DAG) into Dep, and the rest into NonDep. [^2] Note that the answers to all the questions in the Dep set are ‘yes’, and the answers to NonDep questions are ‘no’.
@@
CaT-Bench enables two tasks. Step Order Prediction  elicits binary judgments about dependencies between pairs of steps in a plan, and performance on this task can be evaluated automatically. Step Order Explanation  requires models to provide explanations for their judgments about step dependencies. This involves understanding causal relationships and expressing relevant knowledge about actions in the steps being asked about. Since this is a free-form generation task, these explanations require human evaluation. Note that CaT-Bench does not contain gold, human-written explanations, and we advocate for reference-free human evaluation since there can be multiple valid explanations.
@@ 4.1 Models
We evaluate gpt-4-turbo, gpt-3.5-turbo, gpt-4o, claude-3.5-sonnet, gemini-1.0-pro, gemini-1.5-pro, gemini-1.5-flash, gpt-4o-mini and Llama3-8B. These represent a diverse set from different model families and sizes. We evaluate them primarily in zero-shot prompting modes. We consider two settings: (i) generating only an answer (A), and (ii) generating an explanation along with the answer (A + E). The latter represents answering the question and then generating an explanation for it. We also analyze few-shot results for the answer-only (A) setting, and evaluate generic CoT (E + A) prompting. More details about each model can be found in Appendix A and the prompts used in Appendix D.
@@ 4.2 How Good Are Model Predictions?
Table 1 presents the performance of all the models in different settings on Step Order Prediction. We present per-class (Dep and NonDep) precision, recall and F1 score as well as macro average metrics on the class balanced CaT-Bench. We make three main observations.
@@
In the zero-shot answer-only setting (A), claude-3.5-sonnet records the highest F1 score overall of 0.60. gpt-3.5-turbo and gpt-4-turbo are close behind with 0.59 and 0.58 respectively. Surprisingly gpt-4o, the most recent frontier model, fares significantly worse at 0.49 F1. It’s smaller version, gpt-4o-mini, also performs similarly. All three Gemini models (gpt-3.5-turbo, gemini-1.0-pro, and gemini-1.5-flash) also only manage an F1 of around 0.55. Llama3-8B also fares poorly with an F1 of 0.49. Most models are comparable or barely better than a random baseline F1 of 0.5 on this balanced dataset showing that they are not able to directly answer the dependence question.
@@
Results for adding explanations to answers is shown in the (A + E) rows in Table 1. Seven of the nine models, gpt-3.5-turbo and gemini-1.5-flash being the exceptions, have higher performance when also generating explanations. The biggest improvement in F1 is seen in gpt-4o (+0.21). With explanations, the best result is the 0.73 F1 when using gemini-1.5-pro. While this is substantially better than a random baseline, there is still significant room for improvement.
@@
As noted earlier, using explanations improves the overall performance, translating to more balanced precision/recall values on Dep than when predicting answers alone. Since the bias towards Dep necessarily means bias against NonDep, reduction in bias towards Dep also translates to a more balanced performance on both Dep and NonDep sets. However, even with explanations, the bias towards predicting dependence still remains to some extent for all models. Explanations improve gpt-4o performance the most (+0.36) on NonDep questions. They do not help smaller models (Llama3-8B) identify dependencies better.
@@ 4.3 How Good Are Model Explanations?
Table 2 presents the quality of model generated explanations as judged by human annotators. As expected, larger models are clearly better than the much smaller Llama3-8B on all metrics. There is very little difference between the frontier models, gpt-4o, gpt-4-turbo and gemini-1.5-pro. Avg performance indicates that there is significant room for improvement in the quality of model explanations. On ModAvg, we see that even the best model performance is below 3 (‘neither agree nor disagree’ with a model’s explanation). By this metric, gemini-1.5-pro explanations are worse than GPT-4 even though it generates more correct answers. The difference between Avg and ModAvg indicates models are capable of generating convincing explanations for their wrong answers. They produce explanations which justify the opposite of their answer a significant number of times. In fact, Llama3-8B does so almost half the time. These results show that models have a lot of room for improvement in their ability and reliability to reason about step dependencies in plans.
@@ 5.1 Robustness of Models
Table 3 presents two measures of consistency to quantify the robustness of the models, similar to Verma et al. (2023); Elazar et al. (2021).
@@
For a pair of steps $(step_{i},step_{j})$, the answer to must $step_{i}$ happen before $step_{j}$ should be the same as the answer to must $step_{j}$ happen after $step_{i}$ regardless of question type. As described in subsection 3.1, we measure this notion of consistency through TC. We make two main observations: (i) Even the most consistent models gpt-4o, gemini-1.5-pro and claude-3.5-sonnet change their answers to the before and after versions of questions in 20+of the cases. The rest are far more inconsistent with gemini-1.5-flash changing its answers for more than 55% of the questions; (ii) Surprisingly, adding explanations reduces answer consistency for most models, with gemini-1.5-pro (+24%), gpt-4-turbo (+14%) and claude-3.5-sonnet (+31%) being the only exceptions showing improved consistency upon generating explanations.
@@ 5.2 Chain-of-Thought Struggles
In the experiments thus far, we have asked models to generate explanations for their answers. This can be seen as a answer-then-explain (A + E) approach. In contrast, the standard Chain-of-Thought (CoT) prompting strategy Wei et al. (2022b) can be seen as an explain-then-answer approach (E + A), where we ask the model to generate reasoning or explanation that leads to its answer. In practice, this step-by-step reasoning can be seen as allowing the use of intermediate decoding tokens (like a scratchpad) in service to coming up with a possibly more accurate final answer for many tasks. Table 4 compares performance of CoT prompting [^4] We tried multiple CoT prompts, all with $temp=0$, but it had little effect on performance. (E + A) to first predicting the answer and then explaining it (A + E) and simply providing an answer (A), all in the zero-shot setting when using gpt-4o. [^5] We use gpt-4o instead of gemini-1.5-pro due to rate limits on the latter.
@@
Further manual analysis of the explanations shows that (E + A) generates more vacuous explanations (e.g. just stating that a pair of steps are independent without digging deeper into what makes them independent). Additionally, (E + A) makes errors by misunderstanding entity states for a step. Unsurprisingly, both settings suffer from some hallucinations, sometimes missing the content of the steps being asked about. We present an example of such hallucinations in Figure 4. These results are further indicators of brittleness and inconsistencies in models’ reasoning about step dependencies.
Goal: lightly spiced tomato and red lentil soup Steps:
1. Heat the oil in a saucepan.
2. Add onion and cook over low heat for 7-8 mins… 3. Stir in celery and carrot and cook for 3 mins… 9. Blend the soup using hand-held blender. 10. Check the consistency. 11. Season to taste. Question: Explain why or why not Step 10 must happen after Step 3. Model Answer: Step 10 does not need to happen after Step 3 because the eggplant needs to be cooked before blending. Human Score: 1.0
Figure 4: Example of hallucinations produced by GPT-4 in the (E + A) setting.
We also include zero-shot results with o1-preview, which was released just before the time of publication. o1-preview uses search over chain of thought explanations as part of its inference process. [^6] https://openai.com/index/introducing-openai-o1-preview/ On CaT-Bench, this model achieves state-of-the-art performance scoring 0.80 F1. This is better than both zero- and few-shot performance of any other model, and even achieves the best TC at 85%. However, even this powerful model shows a bias towards predicting dependence (F1 of 0.83) between steps more than their non-dependence (F1 of 0.76). Due to rate limits and prohibitive costs (\$32 for each Step Order Prediction experiment), we were unable to investigate o1-preview further.
@@ 5.3 Effect of Improved Prompting Techniques
We use in-context learning Wei et al. (2022a) with examples selected from the balanced training set using the BM25888https://github.com/dorianbrown/rank_bm25 Robertson and Zaragoza (2009) algorithm. We use k=5 exemplars and dynamically retrieve exemplars from a held-out set that are closest to the test instance. As expected, few-shot prompting improves binary prediction performance a lot (+0.19). In fact, few-shot performance is almost as good as predicting then explaining with gpt-4o.
@@ 6 Conclusion
Understanding plans requires reasoning about its different aspects such as preconditions and effects. This paper introduces CaT-Bench, a new benchmark to evaluate the causal and temporal reasoning abilities about plans. Despite the remarkable strength of current SOTA LLMs, we find that none of them are very good at understanding whether one step in a plan must precede (or succeed) another. Particularly, they are much worse at knowing when there is not a dependency between steps. We also find that LLM predictions are not robust as measured by two metrics of consistency. Prompting LLMs to provide an answer and then to explain it improves performance significantly, and is even better than chain-of-thought (reasoning followed by answering). Human evaluation of these explanations shows that models have a long way to go at understanding dependencies.
@@ Acknowledgements
This material is based on research that is supported in part by the Air Force Research Laboratory (AFRL), DARPA, for the KAIROS program under agreement number FA8750-19-2-1003 and in part by the National Science Foundation under the award IIS #2007290. This material is also based upon work supported by the DARPA’s Perceptually-enabled Task Guidance (PTG) program under Contract No. HR001122C007.